OpenStudy (rational):

help please show that 2^n divides (n+1)(n+2)...(2n)

OpenStudy (tkhunny):

How many terms are in (n+1)(n+2)...(n+n)

OpenStudy (anonymous):

what is the base case?

OpenStudy (anonymous):

does n starts at 0?

OpenStudy (rational):

n terms i think, yes n >= 0

OpenStudy (rational):

sorry it says to prove for n>=1

OpenStudy (anonymous):

ok, use induction. Show me the base case is true

OpenStudy (rational):

okay induction ! n = 1 : 2^1 = 2, (1+1) = 2 2 divides 2 so the statement is true for n=1 assume n=k : 2^k divides (k+1)(k+2)...(2k) show that 2^(k+1) divides (k+2)(k+3)...(2(k+1))

OpenStudy (rational):

lol i have tried everything except induction for the last couple of hours !

OpenStudy (rational):

2^(k+1) = 2*2^k

OpenStudy (anonymous):

yes, the inductive step is always the hardest step.

OpenStudy (anonymous):

So start with 2^(k+1) = 2 * 2^k Do you notice the inductive step is embedded that equation above?

OpenStudy (anonymous):

embedded *in*

OpenStudy (anonymous):

what do we know about 2^k ?

OpenStudy (rational):

2^k divides (k+1)(k+2)...(2k)

OpenStudy (rational):

2^k divides (k+1)(k+2)...(2k) should i multiply 2 both sides ?

OpenStudy (anonymous):

yes but before that, you should write it in terms of equation first. what does 2^k divides (k+1)(k+2)...(2k) mean if expressed in an equation?

OpenStudy (rational):

oh would that be (k+1)(k+2)....(2k) = t*2^k ?

OpenStudy (anonymous):

yes, there exist an integer, and you call it t, so that 2^k (t) = (k+1)(k+2)....(2k) now multiply both sides by 2, what do you get?

OpenStudy (rational):

2^(k+1)t = (k+1)(k+2)...(4k)

OpenStudy (anonymous):

that's mathematical correct, but not helpful

OpenStudy (rational):

i think (k+1)(4k) needs to be written as (2(k+1))

OpenStudy (anonymous):

how about this, 2^(k+1)t = [2(k+1)] (k+2)...(2k) see it?

OpenStudy (rational):

BINGO ! we're done ? xD

OpenStudy (anonymous):

haha yeah, just move the 2(k+1) to the last position 2^(k+1)t = (k+2)...(2k) 2(k+1)

OpenStudy (rational):

Got it completely ! thanks a lot :D

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

notice that (k+2)...(2k) 2(k+1) reall means you start at (k+2) and multiply all the integers in between un till 2(k+1). In other words, (k+2) (k+3) ... 2(k+1)

OpenStudy (rational):

yeaap (k+1 + 1) all the way to 2(k+1) earlier i was trying to factor out 2's from (n+1)(n+2)... (2n) and was hopeless, not sure why it was getting so complicated lol... thank you very much ! induction looks just perfect for this xD

OpenStudy (anonymous):

so u got it ?

OpenStudy (rational):

yes lol got it by induction, still a little curious if we can see it directly

OpenStudy (anonymous):

directly like 2^n | (n+1)(n+2)...(2n) then (n+1)(n+2)...(2n) mod 2^n =0 ?

OpenStudy (rational):

yes exactly ! show that \((n+1)(n+2)...(2n) \equiv 0 \mod 2^n\)

OpenStudy (anonymous):

let me think , i'll grab coffee lol

OpenStudy (rational):

sure :) i got all day... nothing to hurry :)

OpenStudy (rational):

by any chance, you thinking of using WIlson's thm ?

OpenStudy (anonymous):

m back :)

OpenStudy (anonymous):

i was thinking of somthing more dumb than welson lolz

OpenStudy (rational):

im not sure if wilson is useful here lol

OpenStudy (anonymous):

wilson is about if p is prime then (p-1)=-1 mod p well i would think of wilson lol since u mention him ( i wasn't think of this method at all lol )

OpenStudy (anonymous):

how many tems we wud get by add 1 to reach 2n ?

OpenStudy (rational):

Wilson : \((p-1)! \equiv -1 \mod p\)

OpenStudy (anonymous):

yeah made a typo :)

OpenStudy (anonymous):

like for example if n=5 6*7*8*9*10 5 terms .. =n

OpenStudy (rational):

\((n+1)(n+2)(n+3)... (n+n)\) so there are \(n\) terms in the product ?

OpenStudy (anonymous):

(n+1)(n+2)...(2n) what i ment is the terms count is n

OpenStudy (anonymous):

yep n terms :)

OpenStudy (anonymous):

so we wud have 2 cases :- 1- n is odd 2- n is even

OpenStudy (rational):

okay

OpenStudy (anonymous):

im thinking :D

OpenStudy (anonymous):

(n+1)(n+2)...(2n) = (2n)! - n!

OpenStudy (rational):

\((n+1)(n+2)(n+3)... (n+n) \) even : \(n = 2k\) \((2k+1)(2k+2)(2k+3) .... (2k+2k) \equiv (2k)^{n} \equiv 0 \mod 2^n\) odd : \(n = 2k+1\) \((2k+1+1)(2k+1+2)(2k+1+3) .... (2k+1 +2k+1) \equiv (2k)^{n} \equiv 0 \mod 2^n\) QED

OpenStudy (rational):

Oh yes thats get interesting i think : \((n+1)(n+2)(n+3)...(2n) = \dfrac{(2n)!}{n!}\)

OpenStudy (rational):

you meant that right ?

OpenStudy (anonymous):

your methd i think its wrong how would you geranty that 1/2k .... n/2k would be integer the second one yess i ment that since there is a theorm about 2^n

OpenStudy (rational):

\((2k+1)(2k+2)(2k+3) .... (2k+2k) \) multiply all the first terms, you wud get : \((2k)^n\), since you have \(n\) terms right ?

OpenStudy (anonymous):

do you know this theorm ?